OpenStudy (anonymous):

Solve for x. AB DE x = 19 x = 12 x = 9 x = 16

OpenStudy (ajprincess):

I guess that there is a diagram for this. am I right?

OpenStudy (anonymous):

YEA THEIR IS THIS IS IT

OpenStudy (ajprincess):

The ratios of sides \(\triangle DCE\) and \(\triangle ACB\) are equal since similar triangles. That is \(\large\frac{12}{12+x-3}=\large\frac{3}{3+4}\) \(\large\frac{12}{9+x}=\large\frac{3}{7}\) cross multiply and solve for x. Can u do?

OpenStudy (anonymous):

I WILL NEED SOME HELP IN THIS

OpenStudy (ajprincess):

\(\LARGE\frac{12}{9+x}=\LARGE\frac{3}{7}\) Cross multiply. Then u get \(\large7*12=3(9+x)\) \(\large84=27+3x\) Subtract 27 bot sides. Then divide by 3 both sides. Can u do nw?

OpenStudy (anonymous):

SO X =19

OpenStudy (ajprincess):

ya.

OpenStudy (anonymous):

THANKS

OpenStudy (ajprincess):

yw